### v1

@@ 4 Methods
In addition to these general models, specialized GDL models tailored for brain connectome analysis are also evaluated. BrainGB [23] is a systematic study of how to design effective GNNs for brain connectome analysis. The core design dimensions of BrainGB include message passing mechanism and pooling strategies. BrainGNN [3] introduces ROI-aware graph convolutional layers that leverage both topological and functional information from fMRI data. The model also includes ROI-selection pooling layers (R-pool) to highlight the most salient ROIs for interpretability. Regularization terms are used to encourage biologically meaningful ROI selection. NeuroGraph [26] builds on a standard GCN by incorporating a residual connection. It concatenates the hidden embeddings from each GCN layer with the flattened upper triangular part of the brain network matrix. The resulting embedding undergoes batch normalization and is passed through an MLP for final prediction. BrainNetTF [4] applies Transformer-based models to brain connectome analysis, treating each node in the brain network as a ’word in a sentence’. This model utilizes the classic Transformer encoder, which captures pairwise connection strengths among ROIs through self-attention mechanism. Additionally, BrainNetTF introduces an Orthonormal Clustering Readout (OCRead) module, which performs self-supervised soft clustering and orthonormal projection to generate cluster-aware graph embeddings.
@@
All the aforementioned GDL baselines utilize connection profiles as the node features. For the input graphs of each GDL baseline, the original configurations from prior research are adhered to. Specifically, for models such as GCN, GAT, GIN, GraphSage, NeuroGraph, and BrainGNN, only positive correlations between ROIs are retained as edges, based on the common assumption that positive correlations are generally more informative in the context of functional brain networks [3, 26]. In contrast, for BrainGB and BrainNetTF, both positive and negative correlations are incorporated, as outlined in their original papers. The default graph densities for each GDL baseline also follow their original settings, with the default densities set to the top 5% for GCN, GAT, GIN, GraphSage, and NeuroGraph, and the top 10% for BrainGNN. For BrainGB and BrainNetTF, we follow their original settings and use the full graph with 100% density.
@@
Other Related but Not Included Methods: Besides the baselines mentioned above, there are other methods related to our study but are not included as baselines. BrainNetGNN [82] and FBNetGen [25], for instance, incorporate time series as part of node features but adaptively construct brain networks during learning. Since our study focuses on static networks, these methods are excluded. Additionally, recent transformer-based approaches, such as Community-aware BNT [28] and Gated BNT [27], require prior knowledge during training, such as neural system assignments or spatial coordinates for each ROI. As our study uses solely BOLD time-series data and the network matrices derived from them, we do not consider these methods as baselines.
@@
Dual-pathway Model:
The framework of the proposed dual-pathway model is depicted in Fig. 6, which comprises two complementary pathways. The first pathway (orange arrow) retains the simplicity and effectiveness of linear models (LM), using the flattened upper triangular part of the functional brain network matrix as its feature. Formally, for a given sample, let $\mathbf{B}\in\mathbb{R}^{n\times t}$ represent the BOLD time series data, where $n$ is the number of ROIs, and $t$ is the number of time points. The functional brain network is then represented by the Pearson correlation matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, where each element $\mathbf{A}_{ij}$ is the Pearson correlation coefficient between the time series of ROI $i$ and ROI $j$. The feature vector $\mathbf{u}\in\mathbb{R}^{\frac{n(n-1)}{2}}$ is obtained by flattening the upper triangular part of $\mathbf{A}$, which serves as the feature to the LM pathway. This feature construction approach is consistent with the settings in classical ML-based brain connectome analysis studies [17, 12, 18, 19, 20, 21], which enables the model to capture global patterns from the whole connectivity matrices.
The second pathway (blue arrows) captures the graph structure information within the brain connectome. It utilizes the functional brain network as the graph, with node features derived from the BOLD time series embeddings. To obtain these embeddings, the BOLD time series data $\mathbf{B}$ is fed into a 1D-CNN encoder, resulting in node features $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $d$ is the dimension of the time-series embeddings.
@@
$\hat{\mathbf{A}}_{ij}=\begin{cases}\mathbf{A}_{ij},&\text{if }\mathbf{A}_{ij}>0,\\ 0,&\text{otherwise}.\end{cases}$
Following previous works [3, 26], the negative connections are excluded for graph model. This step emphasizes the functional connectivity between ROIs that exhibit co-activation or co-deactivation, reflecting homophily, where the behavior patterns of connected ROIs are similar. This homophily aligns with the core assumptions of commonly used GDLs [83]. In contrast, negative correlations between ROIs, which indicate heterophily or dissimilar behavior patterns, do not fit these assumptions.
@@
  • Performance metrics: For the classification tasks on the ABIDE and PNC datasets, model performance is evaluated using the AUROC score. For the regression tasks on the ABCD and HCP datasets, we use the Pearson correlation coefficient $r$ between the predicted scores and true scores as the performance metric.
@@
  • Training and testing: We randomly split each dataset into three subsets: 70% for training, 10% for validation, and 20% for testing. All deep learning models are trained for 100 epochs using the Adam optimizer with a batch size of 16 and a weight decay of 1e-4. The epoch with the highest AUROC (for classification tasks) or correlation (for regression tasks) on the validation set is used to compare performance on the test set. All reported performances are the average of 10 random runs on the test set with the standard deviation.
@@
  • Parameter tuning: We use the authors’ open-source codes for CPM, CMEP, BrainGB, BrainGNN, NeuroGraph, and BrainNetTF. The graph-based models GCN, GAT, GIN, and GraphSAGE are from the PyTorch Geometric library [84], while the classical ML baselines are the implementation from scikit-learn library [85]. Hyperparameter tuning is conducted via grid search for some important hyper-parameters of these baselines, based on the performance on the validation set. Specifically, for Dual-pathway model, we tune the number of independent training epochs for GAT pathway from {10, 20, 50}, number of layers {2, 3, 4}, and hidden channels {32, 64, 128}; for GCN, we tune the number of layers {2, 3}, hidden channels {32, 128, 256}, and the graph pooling method {concatenation, mean pooling}; for GAT, we adjust the number of layers {2, 3}, hidden channels {32, 128, 256}, and the number of attention heads {2, 4}; for GIN, we vary the number of layers {2, 3}, hidden channels {32, 128, 256}, and $\epsilon$ for neighborhood aggregation {0.0, 0.2, 0.5}; for GraphSAGE, we explore the number of layers {2, 3}, hidden channels {32, 128, 256}, and aggregation method {mean, max}; for BrainGB, we search for the best pooling strategy {concatenation, mean pooling}, hidden dimensions {64, 128, 256}, and message-passing type {weighted sum, node concatenation}; for BrainGNN, we fine-tune the feature dimensions {32, 128} and regularization coefficients {0, 0.1}; for NeuroGraph, the number of layers {2, 3} and hidden channels {32, 64, 128} are optimized; for BrainNetTF, the number of attention heads {2, 4} and clusters {5, 10, 100} are adjusted; for BrainNetCNN, we tune the first and second layer dimensions {16, 32} and {32, 64}, respectively, and the dropout rate {0.3, 0.5}; for all the deep learning models, we also tune the learning rate from {1e-4, 5e-4, 1e-3, 1e-2}; for CPM, the p-value threshold is varied {0.001, 0.01, 0.05, 0.1, 0.5, 1}; for CMEP, the elastic net regularization is fine-tuned with $\alpha$ values {0.01, 0.1, 0.5} and L1 ratios {0.01, 0.05, 0.1, 0.5}; for Logistic Regression, the regularization strength {0.1, 1, 10} and solver {liblinear, lbfgs, saga} are optimized; for ElasticNet, the $\alpha$ values {0.001, 0.01, 0.1, 0.5, 1} and L1 ratios {0.2, 0.5, 0.7} are adjusted; for SVM/SVR, we select between linear and RBF kernels; for Random Forest, the number of trees {100, 200} and maximum depth {10, unlimited} are fine-tuned; and for Kernel Ridge Regression, the $\alpha$ values {0.1, 1, 10} and kernel types {linear, RBF, polynomial} are explored. Additionally, for classical ML models other than CPM and CMEP, the number of the selected features ranges within {100, 200, 500, 1000, 5000, 10000, n(n-1)/2}, where $n$ is the number of ROIs.

### v2

@@ 4 Methods
In addition to these general models, specialized GDL models tailored for brain connectome analysis are also evaluated. BrainGB [braingb] is a systematic study of how to design effective GNNs for brain connectome analysis. The core design dimensions of BrainGB include message passing mechanism and pooling strategies. BrainGNN [braingnn] introduces ROI-aware graph convolutional layers that leverage both topological and functional information from fMRI data. The model also includes ROI-selection pooling layers (R-pool) to highlight the most salient ROIs for interpretability. Regularization terms are used to encourage biologically meaningful ROI selection. NeuroGraph [neurograph] builds on a standard GCN by incorporating a residual connection. It concatenates the hidden embeddings from each GCN layer with the flattened upper triangular part of the brain network matrix. The resulting embedding undergoes batch normalization and is passed through an MLP for final prediction. BrainNetTF [bnt] applies Transformer-based models to brain connectome analysis, treating each node in the brain network as a ‘word in a sentence’. This model utilizes the classic Transformer encoder, which captures pairwise connection strengths among ROIs through self-attention mechanism. Additionally, BrainNetTF introduces an Orthonormal Clustering Readout (OCRead) module, which performs self-supervised soft clustering and orthonormal projection to generate cluster-aware graph embeddings.
All GDL baselines utilize connection profiles as node features, i.e., a row of the Pearson correlation–based functional connectivity matrix. This follows established practice in GDL studies [neurograph, braingb, braingnn, bnt], ensuring methodological alignment with prior work and enabling direct comparability and interpretation within the neuroimaging literature.
@@
For the input graphs of each GDL baseline, the original configurations from prior research are adhered to. Specifically, for models such as GCN, GAT, GIN, GraphSage, NeuroGraph, and BrainGNN, only positive correlations between ROIs are retained as edges, based on the common assumption that positive correlations are generally more informative in the context of functional brain networks [braingnn, neurograph]. In contrast, for BrainGB and BrainNetTF, both positive and negative correlations are incorporated, as outlined in their original papers. The default graph densities for each GDL baseline also follow their original settings, with the default densities set to the top 5% for GCN, GAT, GIN, GraphSage, and NeuroGraph, and the top 10% for BrainGNN. For BrainGB and BrainNetTF, we follow their original settings and use the full graph with 100% density.
@@
Besides the baselines mentioned above, there are other methods related to our study but are not included as baselines. BrainNetGNN [a14030075] and FBNetGen [fbnetgen], for instance, incorporate time series as part of node features but adaptively construct brain networks during learning. Since our study focuses on static networks, these methods are excluded. Additionally, recent transformer-based approaches, such as Community-aware BNT [community_bnt] and Gated BNT [gated], require prior knowledge during training, such as neural system assignments or spatial coordinates for each ROI. As our study uses solely BOLD time-series data and the network matrices derived from them, we do not consider these methods as baselines.
@@
Dual-pathway Model
The framework of the proposed dual-pathway model is depicted in Fig. 7, which comprises two complementary pathways. The first pathway (orange arrow) retains the simplicity and effectiveness of linear models (LM), using the flattened upper triangular part of the functional brain network matrix as its feature. Formally, for a given sample, let $\mathbf{B}\in\mathbb{R}^{n\times t}$ represent the BOLD time series data, where $n$ is the number of ROIs, and $t$ is the number of time points. The functional brain network is then represented by the Pearson correlation matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, where each element $\mathbf{A}_{ij}$ is the Pearson correlation coefficient between the time series of ROI $i$ and ROI $j$. The feature vector $\mathbf{u}\in\mathbb{R}^{\frac{n(n-1)}{2}}$ is obtained by flattening the upper triangular part of $\mathbf{A}$, which serves as the feature to the LM pathway. This feature construction approach is consistent with the settings in classical ML-based brain connectome analysis studies [pos_neg3, pos_neg2, finn2015functional, shen2017using, cui2018effect, PERVAIZ2020116604], which enables the model to capture global patterns from the whole connectivity matrices.
The second pathway (blue arrows) captures the structure information within brain connectome. It utilizes the functional brain network as the graph, with node features derived from the BOLD time series embeddings. This choice of node feature is motivated by our finding that connection profiles, which already encode global relationships among nodes, are highly redundant and thus prone to over-smoothing during message aggregation. In contrast, BOLD time-series embeddings preserve localized and complementary temporal dynamics of each ROI, enabling more effective information propagation across the graph. The localized information contained in the BOLD time-series embeddings also complement the global information modeled in the Linear Modeling (LM) pathway. To obtain these embeddings, the BOLD time series data $\mathbf{B}$ is fed into a 1D-CNN encoder, resulting in node features $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $d$ is the dimension of the time-series embeddings. We choose 1D-CNNs because they are computationally lightweight, less prone to overfitting given the typically limited sample sizes in neuroimaging datasets, and have been widely used in prior studies for modeling fMRI BOLD signals [zhao2025advances].
@@
$\hat{\mathbf{A}}_{ij}=\begin{cases}\mathbf{A}_{ij},&\text{if }\mathbf{A}_{ij}>0,\\ 0,&\text{otherwise}.\end{cases}$ (1)
Following previous works [braingnn, neurograph], the negative connections are excluded for graph model. This step emphasizes the functional connectivity between ROIs that exhibit co-activation or co-deactivation, reflecting homophily, where the behavior patterns of connected ROIs are similar. This homophily aligns with the core assumptions of commonly used GDLs [homophily]. In contrast, negative correlations between ROIs, which indicate heterophily or dissimilar behavior patterns, do not fit these assumptions.
@@
For the classification tasks on the ABIDE and PNC datasets, model performance is evaluated using the AUROC score. For the regression tasks on the ABCD and HCP datasets, we use the Pearson correlation coefficient $r$ between the predicted scores and true scores as the performance metric.
@@
We randomly split each dataset into three subsets: 70% for training, 10% for validation, and 20% for testing. All deep learning models are trained for 100 epochs using Adam optimizer with a batch size of 16 and a weight decay of 1e-4. The epoch with the highest AUROC (for classification tasks) or correlation (for regression tasks) on the validation set is used to compare performance on the test set. All reported performances are the average of 10 random runs on the test set with the standard deviation.
@@
We use the authors’ open-source codes for CPM, CMEP, BrainGB, BrainGNN, NeuroGraph, and BrainNetTF. The graph-based models GCN, GAT, GIN, and GraphSAGE are from the PyTorch Geometric library [pyg], while the classical ML baselines are the implementation from scikit-learn library [scikit_learn]. Hyperparameter tuning is conducted via grid search for some important hyper-parameters of these baselines, based on the performance on the validation set. Specifically, for Dual-pathway model, we tune the number of independent training epochs for GAT pathway from {10, 20, 50}, number of layers {2, 3, 4}, and hidden channels {32, 64, 128}; for GCN, we tune the number of layers {2, 3}, hidden channels {32, 128, 256}, and the graph pooling method {concatenation, mean pooling}; for GAT, we adjust the number of layers {2, 3}, hidden channels {32, 128, 256}, and the number of attention heads {2, 4}; for GIN, we vary the number of layers {2, 3}, hidden channels {32, 128, 256}, and $\epsilon$ for neighborhood aggregation {0.0, 0.2, 0.5}; for GraphSAGE, we explore the number of layers {2, 3}, hidden channels {32, 128, 256}, and aggregation method {mean, max}; for BrainGB, we search for the best pooling strategy {concatenation, mean pooling}, hidden dimensions {64, 128, 256}, and message-passing type {weighted sum, node concatenation}; for BrainGNN, we fine-tune the feature dimensions {32, 128} and regularization coefficients {0, 0.1}; for NeuroGraph, the number of layers {2, 3} and hidden channels {32, 64, 128} are optimized; for BrainNetTF, the number of attention heads {2, 4} and clusters {5, 10, 100} are adjusted; for BrainNetCNN, we tune the first and second layer dimensions {16, 32} and {32, 64}, respectively, and the dropout rate {0.3, 0.5}; for all the deep learning models, we also tune the learning rate from {1e-4, 5e-4, 1e-3, 1e-2}; for CPM, the p-value threshold is varied {0.001, 0.01, 0.05, 0.1, 0.5, 1}; for CMEP, the elastic net regularization is fine-tuned with $\alpha$ values {0.01, 0.1, 0.5} and L1 ratios {0.01, 0.05, 0.1, 0.5}; for Logistic Regression, the regularization strength {0.1, 1, 10} and solver {liblinear, lbfgs, saga} are optimized; for ElasticNet, the $\alpha$ values {0.001, 0.01, 0.1, 0.5, 1} and L1 ratios {0.2, 0.5, 0.7} are adjusted; for SVM/SVR, we select between linear and RBF kernels; for Random Forest, the number of trees {100, 200} and maximum depth {10, unlimited} are fine-tuned; and for Kernel Ridge Regression, the $\alpha$ values {0.1, 1, 10} and kernel types {linear, RBF, polynomial} are explored. Additionally, for classical ML models other than CPM and CMEP, the number of the selected features ranges within {100, 200, 500, 1000, 5000, 10000, n(n-1)/2}, where $n$ is the number of ROIs.
@@
Acknowledgements
This research was supported in part by the US National Science Foundation under Award Numbers 2319449, 2312502, and 2442172, the internal funds and GPU servers provided by the Computer Science Department of Emory University, as well as the University Research Committee of Emory University.